1 solve for x: x/4 -(x+2) / 5 = (2-3x)/ 2
Answer by tommyt3rd(5050)

m20: x/4 -(x+2) / 5 = (2-3x)/ 2
5x-4(x+2)=10(2-3x)
5x-4x-8=20-30x
31x=28
x=28/31